a total of $7000 is invested: part at 9% and the remainder at 11%. How much is invested at each rate if the annual interest is $700
Answer by mananth(16946)

Part I 9.00% per annum ------------- Amount invested =x
Part II 11.00% per annum ------------ Amount invested = y
7000
Interest----- 700.00

Part I 9.00% per annum ---x
Part II 11.00% per annum ---y
Total investment
x + 1 y= 7000 -------------1
Interest on both investments
9.00% x + 11.00% y= 700
Multiply by 100
9 x + 11 y= 70000.00 --------2
Multiply (1) by -9
we get
-9 x -9 y= -63000.00
Add this to (2)
0 x 2 y= 7000
divide by 2
y = 3500
Part I 9.00% $ 3500
Part II 11.00% $ 3500

CHECK
3500 --------- 9.00% ------- 315.00
3500 ------------- 11.00% ------- 385.00
Total -------------------- 700.00

m.ananth@hotmail.ca